# The weakest condition for a Kannan-type contraction mapping on a complete metric space

### v1

@@ Abstract.
In this paper, we investigate the weakest possible conditions for fixed point theorems concerning two types of mappings: Kannan and Chatterjea. We employ the so-called CJM condition, which has been successfully applied to the classical Banach-type mappings by Ćirić [7]. We show that the CJM condition is nearly the weakest possible to have a fixed point, and that a slight modification yields the weakest condition even for these two types of mappings.
@@ 1. Introduction
Several authors [11, 12, 15, 4] have studied generalizations of Kannan mappings, including integral-type conditions and applications in various spaces. In 1974, L. B. Ćirić [7] introduced the generalization of Banach’s contraction principle, which laid the groundwork for what would later become known as CJM contractions. Building on this foundation, J. Matkowski [13] contributed significantly to the development of the CJM framework in 1980 by studying fixed point theorems for contractive mappings in metric spaces, solidifying the theoretical basis for this class of contractions. Finally, J. Jachymski [9] provided a comprehensive study of contractive conditions in 1995, introducing the concept of equivalent conditions for Meir-Keeler-type theorems [14] and further advancing the understanding of generalized contractions.
CJM contractions provide a powerful and unified framework for analyzing fixed point theorems in metric spaces. Specifically, a mapping $T$ on a metric space $(X,d)$ satisfies the CJM condition if the following holds for all $x,y\in X$:
@@
T. Suzuki [17] introduced the Suzuki-type contractions in 2008 which is a generalization that weakens the Banach contraction condition. It not only guarantees fixed points but also characterizes the completeness of the underlying metric space. In 2017, T. Suzuki [18] established significant results using CJM contractions and provided a unified framework for characterizing various contractive conditions, such as Meir-Keeler, Boyd-Wong, and Browder contractions, by using the convergent sequences. However, the article does not explicitly focus on Kannan-type mappings to prove results within the CJM framework. These studies collectively advance our understanding of Kannan-type mappings by developing new analytical techniques and expanding the theoretical scope through generalized contractive conditions see for instance [1, 16, 8].
In this article, Kannan-type mapping has been explored using CJM contractions to generalize the fixed point results. It is achieved using the weakest condition of T.Suzuki’s [18] work to prove fixed point results for Kannan-type mappings as previously not covered by this approach. We also showed the same for Chatterjea-type contraction mapping [6], which is similar to Kannan-type mapping.
@@ 3. The weakest condition for a Kannan-type contraction mapping
First, if $g(n)=\mu$ for all $n\in\mathbb{N}$, we have
@@ 4. In the case of Chatterjea type contraction mapping
First, if $g(n)=\mu$ for all $n\in\mathbb{N}$, we have

### v2

@@ Abstract.
In this paper, we investigate the weakest possible conditions for fixed point theorems concerning two types of mappings: Kannan and Chatterjea. We employ the so-called CJM condition, which has been successfully applied to the classical Banach-type mappings by Ćirić [5]. We show that the CJM condition is nearly the weakest possible to have a fixed point even for these two types of mappings, and that those slight modifications yield the weakest condition.
@@ 1. Introduction
Several authors [1, 3, 6, 9, 10, 13, 14] have studied generalizations of Kannan mappings, including integral-type conditions and applications in various spaces.
Also, Chatterjea [4] showed the existence of a fixed point under a contraction mapping of a Chatterjea type with a structure similar to that of a Kannan-type contraction mapping. A mapping $T:X\to X$ in a metric space $(X,d)$ is called a Chatterjea contraction if there exists a constant $\alpha\in\left[0,\frac{1}{2}\right)$ such that:
$d(Tx,Ty)\leq\alpha\left(d(x,Ty)+d(y,Tx)\right),\text{for all}\ x,y\in X.$
Subrahmanyam [15] showed that Kannan and Chatterjea type contraction mappings, unlike Banach contraction mappings, give a characterization of the completeness of the metric space. In other words, it is equivalent that every Kannan-type contraction mapping on the metric space $(X,d)$ has a fixed point and that the metric space is complete. The same is true for Chatterjea types. In this sense, Kannan and Chatterjea type contraction mappings are natural contraction mappings mathematically.
Furthermore, Kannan-type contraction mappings have a convergence condition based on self-difference, so they are expected to have applications in physics, engineering and other areas of classical methods. In fact, Kannan-type contraction mapping was used in [17] to show the existence of solutions to nonlinear partial differential equations describing damped spring-mass systems and deformations of elastic beams as in engineering. And further, in a point of view in mathematics, graph theory is used and is compatible and works synergistically with Kannan-type contraction mapping. We may expect that Kannan-type contraction mapping may also have applications in data science, which deals with network models depending on the distance between inputs and outputs. However, as this note is a pure mathematics paper, we finish the examples for the applications, and explain the mathematical advances of the theorems.
In 1974, L. B. Ćirić [5] introduced the generalization of Banach’s contraction principle, which laid the groundwork for what would later become known as CJM contractions. Building on this foundation, J. Matkowski [11] contributed significantly to the development of the CJM framework in 1980 by studying fixed point theorems for contractive mappings in metric spaces, solidifying the theoretical basis for this class of contractions. Finally, J. Jachymski [7] provided a comprehensive study of contractive conditions in 1995, introducing the concept of equivalent conditions for Meir-Keeler type theorems [12] and further advancing the understanding of generalized contractions. CJM contractions provide a powerful and unified framework for analyzing fixed point theorems in metric spaces. Specifically, a mapping $T$ on a metric space $(X,d)$ satisfies the CJM condition if the following holds for all $x,y\in X$:
@@
The CJM condition of the Banach-type fixed point theorem is almost the weakest condition, and its modified version, the CJM condition restricted to Picard sequences, is the weakest condition. Suzuki [16] showed that the contraction mapping assuming this weakest condition is equivalent to having a fixed point for Banach type mapping. The contribution of this paper is to introduce affirmative conclusions of Kannan and Chatterjea types to extend the success of Suzuki.
@@ 3. The weakest condition for a Kannan-type contraction mapping
First, if $g(n)=\nu$ for all $n\in\mathbb{N}$, we have
@@ 4. In the case of Chatterjea type contraction mapping
First, if $g(n)=\nu$ for all $n\in\mathbb{N}$, we have